A bus travels 18 km due East at an average speed of 19 m/s, making stops along the way. After that, the bus stops at the bus station for 42 minutes for a lunch break and then continues due East another 22 km for 20 minutes. What is the bus’s average velocity over this entire drive?
1
Expert's answer
2020-12-10T11:04:00-0500
"v=\frac{s_1+s_3}{\frac{s_1}{v_1}+t_2+t_3}\\
v=\frac{18+22}{\frac{18\cdot60}{19000}+42+20}=0.645\frac{km}{min}=38.7\frac{km}{h}"